OpenStudy (anonymous):

find the reference angle 9╥/4 in degrees

OpenStudy (anonymous):

Do you know the conversion formula?

OpenStudy (anonymous):

yeah im just confused if its the angle toward the x axis or y axis? so it would be 1/4 but which side?

OpenStudy (anonymous):

It's over 2pi, so I'll give you a hint, it will be over 180^o

OpenStudy (anonymous):

its over 360 I know

OpenStudy (anonymous):

so I use the angle connecting the 1/4 radiant and y axis??

OpenStudy (dumbcow):

no "x-axis" for reference angle the reference angle or terminal angle is always the smallest angle from the x-axis \[\frac{9\pi}{4} = \frac{8\pi}{4} +\frac{\pi}{4} = 2\pi + \frac{\pi}{4}\]